Given the linear transformation below:

T (x1, x2, x3) → (x1-x2+2x3, 2x1-2x3, -x1-x2+4x3, 3x1-x2)

T = R3 → R4

1. Determine the transformation matrix of the linear transformation above

2. Determine Ker(T) and Range(T)


1
Expert's answer
2021-11-29T13:21:35-0500

1.

"Ax=\begin{pmatrix}
 1 & -1&2 \\
 2&0 & -2\\
-1&-1&4\\
3&-1&0
\end{pmatrix}\begin{pmatrix}
 x_1 \\
 x_2 \\
x_3\\

\end{pmatrix}=\begin{pmatrix}
 x_1-x_2+2x_3 \\
 2x_1-2x_3 \\
-x_1-x_2+4x_3\\
3x_1-x_2
\end{pmatrix}"


transformation matrix:

"A=\begin{pmatrix}
 1 & -1&2 \\
 2&0 & -2\\
-1&-1&4\\
3&-1&0
\end{pmatrix}"


2.

kernel:

"x_1-x_2+2x_3=0"

"2x_1-2x_3=0"

"-x_1-x_2+4x_3=0"

"3x_1-x_2=0"


"x_1=x_3,x_2=3x_1"

"ker\ T=span(1,3,1)"


range:

"range\ T=\begin{pmatrix}
 x_1-x_2+2x_3 \\
 2x_1-2x_3 \\
-x_1-x_2+4x_3\\
3x_1-x_2
\end{pmatrix}=x_1\begin{pmatrix}
 1 \\
 2 \\
-1\\
3
\end{pmatrix}+x_2\begin{pmatrix}
 -1 \\
 0 \\
-1\\
-1
\end{pmatrix}+x_3\begin{pmatrix}
 2 \\
 -2 \\
4\\
0
\end{pmatrix}="


"=span((1,2,-1,3),(-1,0,-1,-1),(2,-2,4,0))"